OpenStudy (anonymous):

does anyone know how to write 57/11 as mixed number

OpenStudy (anonymous):

57/11 = 55/11 +2/11 = 5 +2/11

OpenStudy (anonymous):

so the answer is 5 2/11?

OpenStudy (anonymous):

yes :D

OpenStudy (anonymous):

thanks can you help with anotehr

OpenStudy (anonymous):

shoot it over :D

OpenStudy (anonymous):

use order of operatiorn to simplify 15+3 over 15-3

OpenStudy (anonymous):

I need to see the problem, because the answer could either be 3/2 or it could be 12+1/5

OpenStudy (anonymous):

mmmm well it is written 15 +3 and the line under then 15 -3

OpenStudy (gw2011):

If you mean (15+3)/(15-3), then this becomes: 18/12 = 3/2 = 1 1/2